OpenStudy (anonymous):

Write each fraction as a percent. A. 7/20 B. 10/50 C. What method did you use to do this?

jigglypuff314 (jigglypuff314):

convert the fractions into decimals then multiply the decimals by 100 to get percent for example like fraction 1/5 = 0.20 as a decimal = 20%

OpenStudy (anonymous):

Divide the top number (numerator) by the bottom number (denominator). For example, on Question A: 7 divided by 20 = 0.35 = 35% 0.35 = 35% because to turn a decimal into a percent, you move the decimal place two places to the right.

OpenStudy (grace111):

a) 7/20 - 7 divided by 20 =0.35 and 0.35 x 100 = 35% b) 10 / 50 - 1 divided by 5 = 0.2 and 0.2 x100 = 20% Thats how my friend tought me hope this has helped :)